OpenStudy (bbb911):

This is Geometry: Find the volume, to the nearest hundredth, of a sphere with a radius of 3 meters.

OpenStudy (amistre64):

\[V_{sph}=\frac{4}{3}pi r^3\]

OpenStudy (anonymous):

\[\frac{4}{3}\times \pi \times 3^3\]

OpenStudy (amistre64):

3^3 = 27 .... 27/3 = 9 ...soo 9*4pi = 36pi

OpenStudy (bbb911):

if its rounded would it be 37.70 m cube

OpenStudy (amistre64):

\[v=4\pi\frac{3^3}{3}\] is another way to write it \[v = 4\pi\frac{3.3.3}{3}=4\pi3(3)=4\pi9=36pi\]

OpenStudy (amistre64):

let me check...

OpenStudy (amistre64):

113.097.... is abt 36pi

OpenStudy (amistre64):

so to round its; 113.10

OpenStudy (bbb911):

ooo thanks so much :D

OpenStudy (amistre64):

yw :)